OpenStudy (hy123):

C8H18(g)+O2(g) ---> CO2(g)+H2O(g) How many moles of water are produced in this reaction? Hint: Take the number of moles of oxygen and divide it by the coefficient for oxygen. Then multiply by the coefficient for water.

OpenStudy (anonymous):

first start by balancing reaction, then you need to have some starting data mass or volume of reactants etc. then you take the ratio from the balanced equation like this: CH4 + 2O2 -> CO2 + 2H2O n(H2O)/n(CH4)=2/1 = 2 so: n(H2O)= 2 n(CH4) or n(H2O)/n(O2) = 2/2 = 1 so: n(H2O) = n(O2)